OpenStudy (anonymous):

Gravity and Electromagnetic force Compare and contrast these forces. Also provide two ways that they are alike and two ways that they are different.

OpenStudy (anonymous):

Both of these forces are fundamental forces of nature, as in one cannot explain the other. Both of these forces are also interact in the similar expression of inverse square law of interaction; refer to the formulae given, F varies with 1/r^2. Electromagnetic forces are repulsive/ attractive in nature depending upon the charges. The have two different types of interacting fields: Electric and Magnetic. The change in one of these fields creates the other, so their nature is dynamic. Gravitational force, on the other hand, is always attractive. Also it has only one kind of interacting field. Gravity has a static field; it only varies with distance but the variation does not create any interaction apart from classical force of gravity. Electromagnetic force is given by Coulomb's law (for two charges)\[F_E =k\frac{q_1q_2}{r^2}\]and the gravitational force is given by Newton's law (for two masses)\[F_G=G\frac{M_1M_2}{r^2}\]Consider these two forces for two electrons and find their ratio \[\frac{F_E}{F_G}\] This will give you an accurate idea of their relative strengths.